OpenStudy (anonymous):

Find the length of the missing side. leave your answer in simplest radical form

OpenStudy (anonymous):

\[a)\sqrt{29m} \] \[b)\sqrt{445m}\] \[c) \sqrt{7m}\] \[d) \sqrt{203m}\]

OpenStudy (nikato):

Can you try setting up the equation on your own?

OpenStudy (anonymous):

18^2 =324 11^2= 121 324+121=445 so b ?

OpenStudy (nikato):

The first 2 are correct. But you're actually subtracting. 324-121 11^2 + b^2 = 18^2 121 + b^2 =324 Now try to solve for b

OpenStudy (anonymous):

it is D

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c ^{2}\]